a square deck has a perimeter of 50 ft.what is the approximate length of the diagonal of the square deck?
Answer by checkley77(12844)

50/4=12.5 FT. IS THE LENGTH OF THE SIDES OF THE SQUARE DECK.
A^2+B^2=C^2
12.5^2+12.5^2=C^2
156.25+156.25=C^2
C^2=312.5
C=SQRT312.5
C=16.68 FT. IS THE LENGTH OF THE DIAGONAL.